OpenStudy (anonymous):

The dimenson of Triangle B are 3 times the dimensions of Triangle A the area of trinagke B is 72cm2........What is area of triangle A

OpenStudy (anonymous):

\[(l _{1}/l _{2})^{2} = Area _{1} / Area _{2}\]You can take any length to be length one but you must take its area as area one. I hope this helps.

Directrix (directrix):

If two polygons are similar, then the square of the scale factor is the ratio of the areas of the two triangles. (1/3)^2 = A/72 where A is the area of the smaller triangle 1/9 = A/72 9A = 72 A = 8 square cm

OpenStudy (anonymous):

Thank you so much